0 0
Add a comment Improve this question Transcribed image text
Answer #1

a) taking the equations of linear motion into consideration,

s = ut + \frac{1}{2}a t^{2}

0.028 = 0 + \frac{1}{2}a (2.1\times 10^{-8})^{2}

a = 1.27\times 10^{14}m/s^{2}

now. using this acceleration,

v = u + at

v = 0+ at

v = 1.27\times 10^{14}\times 2.1\times 10^{-8}

v = 2.67\times 10^{6}m/s

b) Force acting on the elctron is,

ma = qE

9.1\times 10^{-31}\times 1.27\times 10^{14} = 1.6\times 10^{-19}E

E = 722 N/C